OpenStudy (outnot):

Anyone mind helping me understand this question? if f(x)=5x-4 and g(x)=x^2, what is the (g o f)(2)

OpenStudy (hacer):

g(f(2))= F(2)=5.2-4=6 and g(6)=6^2=36

OpenStudy (outnot):

oh thank you!

OpenStudy (outnot):

how do you give medals?

OpenStudy (hacer):

You have a symbol about medals on your answer or the other users

OpenStudy (outnot):

thanks again

OpenStudy (hacer):

😊

OpenStudy (outnot):

mind helping with another?

OpenStudy (outnot):

im trying to finish this test so i can just enjoy my summer haha

OpenStudy (hacer):

If I can sure

OpenStudy (outnot):

ok thanks one sec

OpenStudy (hacer):

Ok

OpenStudy (hacer):

If I can sure

OpenStudy (outnot):

if f(x)=-x+8 and g(x)=x^4, what is the (f o g)(2)

OpenStudy (hacer):

f(g(2))= g(2)=-2+8=6 f(6)=6^4=1296

OpenStudy (outnot):

thanks again

OpenStudy (hacer):

you are welcome

OpenStudy (outnot):

so are you just taking the 2 and multiplying?

OpenStudy (hacer):

in your second question first you must put 2 instead of x your function g(x) then when you find g(x) put f(g(x)) because it is compound function

OpenStudy (outnot):

ah ok

OpenStudy (outnot):

do you want to keep going or are you gonna move on?

OpenStudy (outnot):

sorry im really new to how the people of open study are.

OpenStudy (hacer):

ok

OpenStudy (hacer):

when people answer your question and you understand you should close question

OpenStudy (outnot):

oh ok thanks so ill do that haha

OpenStudy (hacer):

ok